OpenStudy (anonymous):

Will become a fan and give a medal. Which expression is a perfect cube? answers below. @usukidoll

OpenStudy (anonymous):

D or B

OpenStudy (usukidoll):

mmm... perfect cube numbers are like 8 because 2^3 = 8 then there's 27 which is 3^3 = 27 then 64 4^3 = 64

OpenStudy (usukidoll):

there's more than 1 in here.

OpenStudy (anonymous):

ok

OpenStudy (usukidoll):

ah I see.. for a perfect cube (for exponents) we have to divide by 3

OpenStudy (usukidoll):

like for example